Session: Nov 1999 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Sergio Mariotti at Fri Nov 19 10:44:20 MET 1999) Minor failures
NO scans lost. Sometime, few seconds pointing off due short blackout of Main Electrical Power. Weather rain, snow, wind. RFI on band

Torun (K.Borkowski at Fri Nov 19 10:54:57 MET 1999) Minor failures
No problems noted until 6:05 of day 323. Then, a power failure caused a damage to a few sensors of the elevation drive resulting in a loss of telescope control. The observations could not be continued. Work is in progress to purchase new sensors and bring them from Warsaw yet today, so that the telescope can be made operational before the next experiment in this session.

Jodrell (Alastair Gunn at Fri Nov 19 12:35:10 MET 1999) Minor failures
Telescope servo failed at 03:57 UT and restored at 04:17 UT. Brief servo failures also at 04:47 and 04:56 UT. Azimuth motor failure at 08:26 UT, restored by 09:05 UT. No other known problems.

Cambridge (Alastair Gunn at Fri Nov 19 12:35:46 MET 1999) Minor failures
Telescope control failure at 08:26 UT, restored by 09:05 UT. No other known problems.

Effelsberg (A. Kraus at Fri Nov 19 12:36:02 MET 1999) Failure
A significant number of scans had to be skipped due to snowfall. Observations were only possible between UT2200-0030 and UT0600-0630.

Westerbork (Tony Foley at Mon Nov 22 09:00:16 MET 1999) Success
No known problems. 13 telescopes in the tied array.

Onsala (A. Polatidis at Mon Nov 22 10:16:01 MET 1999) Success

Noto (C.Trigilio at Thu Nov 25 10:52:35 MET 1999) Success
Only 4 minuts lost because of power fail.

Calibration (Jean\-Francois Desmurs at Wed Dec 15 13:13:09 MET 1999) Minor failures
The ANTAB formatted file for this experiment can be found on VLBEER in the nov99 directory under the name: ep032.antab. Read the comments in the file to get further information. (Rem : Cm no Tsys values (all values are nul))